OpenStudy (anonymous):

Differentiate the function q(x) = 4sqrt(x+3)/(x-1)

OpenStudy (anonymous):

use the quotient rule then use the chain rule for when you have to differentiate the sqrt(x+3)

OpenStudy (anonymous):

\[q(x)=\sqrt[4]{(x+3)/(x-1)}\] is this the equation that your answer was solving?

OpenStudy (anonymous):

you can use u-substitutions: u=(x+3)/(x-1) v=\[u ^{1/4}\] u'*v'= your answer u'= -4/(x-1)^2 v'=(1/4)*v^(-3/4) = (-1)/ ((x-2)^2*((x+3)/(x-1))^3/4)

OpenStudy (anonymous):

\[4\times \sqrt{x+3}/(x-1) \] This is how I understood the q.